# Whisper D-SGD: Correlated Noise Across Agents for Differentially Private Decentralized Learning

### v1

@@ VI-A2 Algorithms
We compare (i) LDP [13, 14], which adds independent local noise; (ii) DECOR [18], representative of pairwise correlation; and (iii) Whisper D-SGD, our mixing-based correlation approach. All algorithms run for $T=5000$ iterations. Results are averaged over 10 random seeds.
@@ VI-A3 Tasks
We consider two machine learning tasks:
@@ VI-B Experimental Results
Figures 33 show the empirical privacy-utility trade-off on both (i) the quadratic optimization task (reporting optimality gap $\frac{1}{n}\!\sum_{i=1}^{n}f_{i}(x_{i}^{(T)})-f^{*}$) and (ii) the real-world logistic regression task (reporting test loss). Panel (a) of each figure varies the privacy budget $\varepsilon$ under fixed connectivity $p=0.5$, while panel (b) varies $p$ under fixed $\varepsilon=10$.
Across all $\varepsilon$ values, both correlation-based approaches (DECOR and Whisper D-SGD) significantly outperform LDP, decreasing final losses by up to an order of magnitude (Figs. 2(a)2(c)). This confirms that correlated noise effectively mitigates the performance drop caused by local noise.
Whisper D-SGD consistently outperforms DECOR for all network connectivities (Figs. 2(b)2(d)). For fully connected networks ($p=1$), $\mathbf{W}$ and $\mathbf{L}$ become structurally equivalent, therefore both methods achieve nearly the same utility. However, at lower connectivities ($p\leq 0.4$), Whisper D-SGD outperforms DECOR by the largest margin, confirming the importance of $\mathbf{W}$ in the correlated-noise design.
In summary, Whisper D-SGD significantly enhances the utility–privacy trade-off, particularly in sparse or moderately connected networks. Our results demonstrate the benefits of mixing-aware correlation compared to baseline approaches that disregard network connectivity and mixing weights.
@@ VII Conclusion
In decentralized learning systems, sensitive information about training data can be exposed through shared local models. While LDP mitigates such privacy risks, it suffers from reduced utility compared to CDP due to accumulated local privacy noise. To address this challenge, we propose the Whisper D-SGD algorithm, which generates correlated noise across agents using a covariance-based approach. This method incorporates the mixing matrix into the design of correlated noise, providing strong privacy guarantees with minimal impact on utility. Both theoretical analysis and empirical results demonstrate that Whisper D-SGD improves the privacy–utility trade-off compared to existing methods (e.g., LDP, pairwise correlation) under various privacy budgets, particularly in sparse network topologies. Future directions include extending this framework to directed graphs and jointly optimizing mixing weights and the covariance matrix to improve convergence and further enhance noise cancellation.

### v2

@@ VI-A2 Algorithms
We compare (i) LDP [13, 14], which adds independent local noise; (ii) DECOR [18], representative of pairwise correlation; and (iii) CorN-DSGD, our covariance-based approach. All algorithms run for $T=5000$ iterations. Results are averaged over 10 random seeds.
@@ VI-A3 Tasks
We consider three machine learning tasks:
@@
(iii) Image Classification. We train a shallow neural network on the MNIST dataset [24], with the same train-test split, data partitioning scheme, and hyperparameter grid as in task (ii).
@@ VI-B Experimental Results
Figures 45 show the empirical privacy-utility trade-off on (i) the quadratic optimization task (reporting optimality gap $\frac{1}{n}\!\sum_{i=1}^{n}f_{i}(x_{i}^{(T)})-f^{*}$), (ii) logistic regression task (reporting test loss), and (iii) image classification task (reporting test accuracy). Panel (a) of each figure varies the privacy budget $\varepsilon$ under fixed connectivity $p=0.5$, while panel (b) varies $p$ under fixed $\varepsilon=10$. Across all $\varepsilon$ values, both correlation-based approaches (DECOR and CorN-DSGD) outperform LDP (Figs. 45(a)). CorN-DSGD consistently outperforms DECOR for all network connectivities (Figs. 45(b)). For fully connected networks ($p=1$), $\mathbf{R}_{\text{cor}}$ and $\mathbf{L}$ become proportional, therefore both methods achieve nearly the same utility. However, at lower connectivities ($p\leq 0.4$), CorN-DSGD outperforms DECOR by the largest margin, confirming the benefits of our covariance-based framework compared to prior approaches in sparse networks.
@@ VII Conclusion
In serverless decentralized learning, sensitive information about training data can be exposed through shared local models. While LDP mechanisms mitigate such privacy threats, they suffer from reduced utility due to independently added local privacy noise. We propose the CorN-DSGD algorithm, which generates correlated noise across agents using a covariance-based approach. This method incorporates the mixing weights into the design of correlated noise, offering strong privacy guarantees with minimal impact on utility. Both theoretical analysis and empirical results demonstrate that CorN-DSGD improves the privacy-utility trade-off compared to existing methods (LDP, DECOR) under various privacy budgets, particularly in sparse network topologies.